OpenStudy (anonymous):

PROBLEM SOLVING USING SYSTEMS OF EQUATIONS There are 28 coins that are all nickles and dimes. The value is $2.05. How many nickles are there and how many dimes?

OpenStudy (lgbasallote):

let x = no. of nickels y = no. of dimes x + y = 28 5(cents)x + 10(cents)y = 205 cents multiply the first equation by -10 -10x -10y = -280 5x + 10y = 205 -5x = -75 x = 15 There are 15 nickels and 13 dimes.

OpenStudy (precal):

you can also use x+y=28 .05x+.10y=2.05

OpenStudy (anonymous):

N + D = 28 5N + 10D = 205 -5(N + D = 28) 5N + 10D = 205 -5N - 5D = 140 5N + 10D = 205 -------------------------- 5D = 65 D = 13 N = 15 There were 13 dimes and 15 nickels

OpenStudy (anonymous):

thankyou so much!

OpenStudy (anonymous):

no problem